Jefferson Lab Weekly Briefs June 22, 2011
Jefferson Lab Weekly Briefs
June 22, 2011
12 GeV Upgrade
Work is progressing on the five superconducting magnets for the Hall C Super High Momentum Spectrometer (SHMS). Materials orders are being placed for the horizontal bend (HB) dipole magnet, and manufacturing design reviews are scheduled for the Q1 quadrupole magnet. Initial samples of superconducting cable soldered to the extruded copper stabilizer have been prepared and sent to the vendor for the SHMS Q2 and Q3 quadrupole and the dipole magnets. Bids have been received for the wheels and bogies supporting the main spectrometer carriage, and the requests for proposals are being readied for the various large sections of the support carriage.
During the week, Hall A continued the installation of the target and associated equipment for the upcoming g2p/gep experiments. The installation is currently on schedule. There was a special meeting to discuss how to determine the transport optics for these experiments. This was productive and no additional equipment is going to be needed; although experiment collaborators were exploring the idea of using a thick sieve plate instead of the planned thin one, providing a suitable thick plate can be found. There was discussion of requesting half-pass running during the Hall A special energy period; but no decision has been made at this point.
The fifth week of the six month shutdown went well. The beamline valve interlock and solenoid were installed for the injector project (PEPPo). Radiofrequency cryomodule C100-1 was moved from the tunnel to the Test Lab for more measurements and possible improvement. Installation of the west arc dipoles continued; about one-third of the dipoles were installed and beamline connection and alignment were in progress. Installation of RF infrastructure (including cabling, control hardware, MO distribution, waveguides, klystrons and etc.) for the new zones in the south linac had been going well. Many other tasks were completed, continued or started as scheduled.
The FEL team warmed up the booster cryounit last week to room temperature and began the staging of replacement parts in the injector.
Center for Theoretical and Computational Physics
One of the main physics goals of a future Electron Ion Collider (EIC) will be to accurately measure nuclear modifications of gluon and quark parton distribution functions (PDFs), as well as the possible onset of nonlinear QCD dynamics (saturation) in heavy nuclei. By performing global QCD fits to "pseudo data" for inclusive deep inelastic scattering cross sections on a lead target at an EIC, it is shown that nuclear PDFs can be obtained with a precision comparable to current global fits, but over an extended kinematic range (e.g., for gluon shadowing down to momentum fraction x = 10-3), and the gluon anti-shadowing and EMC effects can be measured at x > 0.1. Furthermore, saturation signals may already be detected at a medium-energy EIC (√s = 12-44 GeV).
The RF (EESRF) Group has been engaged in the installation of 12 GeV RF equipment in south linac zones 24 and 25. The cable pulls to the tunnel for cryomodule control have been completed. Mechanical and electrical sub-assemblies are being installed in the equipment gallery. Items being placed include waveguide components, water cooling manifolds, klystron cabinets, klystron cathode power supplies and support electronics, and low-level RF control chassis and cabling interconnections. Klystron installation should begin this week. A temporary water cooling system has been assembled and placed behind the south linac building to support commissioning of the hardware while the accelerator low-conductivity water system undergoes upgrades. Initial testing of the first RF zone is schedule to begin in a few weeks.
The instrument and controls (EESIC) group has been engaged in the CAMAC to VME controls upgrade in buildings W2 and W4. Additionally, the group is also working on the SEE beam position monitoring electronics for Arc 10 in the west arc and final production testing and installation of the SRF vacuum controls.
The safety system (EESSAF) group has been modifying the south linac personnel safety system to support the new RF zones.
The magnet power (EESDC) group has been preparing and testing two power supplies to be installed in Hall A for the g2p experiment. The group is also installing cables for new trim power supplies in the west arc and box power supply infrastructure upgrades. Both items are 12 GeV projects.
Support group electricians (EESES) have been installing cable trays and cabling for the power supply reconfigurations planned for the 12 GeV Upgrade. Electricians are also hooking up the electrical connections for the 12 GeV refurbished magnets as they are placed back in the tunnel. The fabrication shop is busy building and assembling items in support of all these ongoing activities.
The electric engineering CAD (EECAD) team is tasked with supporting both new construction documentation as well as capturing all of the changes being implemented during the shutdown.
Environment, Safety, Health and Quality
Many injuries occur due to items on the floor that cause slips, trips or entanglements. These injuries can be easily prevented with a little housekeeping.
It is up to all of us to help provide a safe work environment for us and our fellow employees. If you add clutter to your work environment during the course of your duties, make sure you leave time to clean up your mess as soon as practical after making it. Never leave items on the floor, which could cause a slip, trip or fall, and never expect others to clean up your mess.
New Montage Posted
JLab Initiatives Fund Call for Proposals Announced
Smoke Shacks Gain Ground
Semi-Annual Personal Dosimeter Change Out Set for June 29
Minority Undergraduate Research Assistantship
Tracking Thomas at JLab
Congratulations to Jason Willoughby, who was the first person to correctly identify the June 15 location. Honorable mentions go to Doug Higinbotham, Marcy Stutzman, Michael Spata, Marie Ivanco, Karl Slifer, Andre Newman, Stephen Smith, Dick Owen, Harry Fanning, David Fazenbaker, Stephanie Vermeire, Ron Bartek, Mike Davis and Mike Koch. Check out the Tracking Thomas webpage for a better view of his last location and this week's new mystery photo.
JLab Calendar of Events
July 4: Independence Day holiday; JLab closed
JLab Summer 2011 Education Calendar
May 23 - July 29: Science Undergraduate Laboratory Internship